OpenStudy (anonymous):

Two trains running in opposite directions cross a man standing on the platform in 27 seconds and 17 seconds respectively and they cross each other in 23 seconds. The ratio of their speeds is?

OpenStudy (anonymous):

\[Speed_{1} = \frac{d_1}{t_1 } = \frac{d_1}{27}\]

OpenStudy (anonymous):

\[Speed_2 = \frac{d_2}{t_2} = \frac{d_2}{17}\]

OpenStudy (anonymous):

When trains will be running in opposite, their relative speed will add up: So: \(Speed = Sp_1 + Sp_2\)

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

then?

OpenStudy (anonymous):

@eta ?